# RadMamba: Efficient Human Activity Recognition through Radar-based Micro-Doppler-Oriented Mamba State-Space Model

### v1

@@ Abstract
Radar-based Human Activity Recognition (HAR) has emerged as a promising alternative to conventional monitoring approaches, such as wearable devices and camera-based systems, due to its unique privacy preservation and robustness advantages. However, existing solutions based on convolutional and recurrent neural networks, although effective, are computationally demanding during deployment. This limits their applicability in scenarios with constrained resources or those requiring multiple sensors. Advanced architectures, such as Vision Transformer (ViT) and State Space Model (SSM) architectures, offer improved modeling capabilities and have made efforts toward lightweight designs. However, their computational complexity remains relatively high. To leverage the strengths of transformer architectures while simultaneously enhancing accuracy and reducing computational complexity, this paper introduces RadMamba, a parameter-efficient, radar micro-Doppler-oriented Mamba SSM specifically tailored for radar-based HAR. Across three diverse datasets, RadMamba matches the top-performing previous model’s 99.8% classification accuracy on Dataset DIAT with only 1/400 of its parameters and equals the leading models’ 92.0% accuracy on Dataset CI4R with merely 1/10 of their parameters. In scenarios with continuous sequences of actions evaluated on Dataset UoG2020, RadMamba surpasses other models with significantly higher parameter counts by at least 3%, achieving this with only 6.7 k parameters. Our code is available at: https://github.com/lab-emi/AIRHAR.
Index Terms:
Human activity recognition, Mamba state-space model, Frequency Modulated Continuous Wave (FMCW) radar, micro-Doppler signatures, deep learning, parameter-efficient networks, continuous monitoring, radar signal processing
@@ I Introduction
Human Activity Recognition (HAR) technologies have become increasingly critical in modern society, finding essential applications in healthcare, elderly care, smart homes, and security. Traditional HAR methods predominantly rely on wearable sensors or camera-based systems. While wearable devices offer precise monitoring, they often compromise user comfort and face challenges related to battery life and user compliance. Camera-based solutions, although versatile, introduce significant privacy concerns and are susceptible to performance degradation under challenging environmental conditions. Consequently, Radar-based Human Activity Recognition (RadHAR) systems have emerged as a compelling alternative due to their inherent advantages, including user privacy preservation, robustness to light conditions, non-intrusive and through-wall sensing capabilities [1, 2].
The field of radar-based HAR has evolved considerably from initial works focusing on simple scenarios with clearly defined activities [3, 4] toward more complex and realistic settings, including continuous activity monitoring [5, 6], multi-modal and distributed sensing [7, 8], and advanced classification techniques based on machine learning and deep learning methods. Early solutions predominantly relied on conventional machine learning algorithms, such as Support Vector Machines and k-Nearest Neighbors (kNN), utilizing handcrafted features like micro-Doppler centroid, bandwidth, and statistical measures [4, 5, 6]. Although these traditional techniques demonstrated acceptable performance in constrained scenarios, they encountered challenges related to scalability, complexity in feature engineering, and generalization to realistic continuous environments.
The advent of deep learning techniques, particularly convolutional neural networks (CNNs) [9, 10], recurrent neural networks (RNNs) [11, 7], and hybrid CNN-RNN models [12, 13, 14], revolutionized radar-based HAR by enabling automatic extraction and representation of complex spatial-temporal features directly from radar data. However, despite their accuracy, CNN and hybrid models typically suffer from high computational complexity and latency, limiting their suitability for real-time deployment, especially in resource-constrained edge scenarios.
Recently, advanced transformer-based architectures such as ViT and SSM have demonstrated significant improvements in feature modeling capabilities and explored lightweight designs across various domains [15, 16, 17, 18, 19, 20]. Nevertheless, their computational complexity remains high, making it challenging to balance the number of parameters and the number of floating-point operations per inference (#FLOP/Inf.). For instance, the FML-ViT model with only 169 M #FLOP/Inf. requires 2.7 M parameters [16], while model LW-ViT with 769 k parameters demands significantly higher #FLOP [21]. Compared to ultra-lightweight object detection models optimized for edge devices, such as YOLO-Fastest with 350 k parameters, and 252 M #FLOP/Inf, the direct application of these advanced architectures to radar-based HAR tasks remains costly due to their high computational overhead. This limitation is particularly critical for applications with limited resources, such as gesture recognition in assistive technologies on mobile or embedded devices [22]. Additionally, applications like smart home monitoring, which often involve multiple sensors deployed across a home, would greatly benefit from lightweight and energy-efficient algorithms [8].
To overcome these limitations and leverage the modeling strengths of state-space architectures while ensuring computational efficiency in radar-based HAR, we introduce a novel SSM Mamba-based architecture named RadMamba. In summary, the main contributions of this work are as follows:
  • We propose RadMamba, a parameter-efficient radar-based HAR architecture incorporating Channel Fusion and Downsampling, Doppler-aligned segmentation, and convolutional token projections, specifically tailored to address the unique characteristics of micro-Doppler radar signatures.
  • RadMamba achieves state-of-the-art classification accuracy on multiple radar-based HAR datasets: it matches the best accuracy of existing heavyweight models on non-continuous datasets (99.8% accuracy on Dataset DIAT and 91.2% accuracy on Dataset CI4R), and surpasses existing models by at least 3% on scenarios with continuous sequences of activities (89.3% accuracy on Dataset UoG2020), all while significantly reducing the number of parameters. Moreover, its computational complexity is the lowest compared to previous models validated on Dataset CI4R and Dataset UoG2020.
  • We perform a comprehensive ablation study demonstrating the critical role and synergy of each proposed architectural component, showing that the combination of micro-Doppler-aligned segmentation, adaptive downsampling, and convolutional projections is essential for optimal performance.
@@
  • We provide an open-source PyTorch framework, RadHAR, facilitating rapid prototyping and evaluation of micro-Doppler-based HAR methods, promoting reproducibility and future research.
@@ II-A Conventional machine learning method
These methods involved extracting domain-specific features from radar signals, followed by classification using algorithms like kNN and SVM. For instance, Erol et al. [5] used kNN to analyze power burst curves in spectrograms for fall detection, distinguishing fall from non-fall events. Li et al. [6] applied SVMs to micro-Doppler features such as Doppler centroid, bandwidth, and Singular Value Decomposition (SVD) components, while Ding et al. [23] extracted 28 features of four types: dynamic Doppler frequency, range change, energy change and dispersion of range and Doppler from dynamic range-Doppler trajectories for subspace kNN classification. Although effective in controlled environments, these approaches struggled to generalize across diverse conditions and activities due to their dependence on hand-crafted features, often leading to overfitting and poor scalability for continuous tasks [1, 2].
@@ II-B CNN-based method
The shift to deep learning marked a turning point for RadHAR, with Convolutional Neural Networks leading the charge. Unlike traditional methods, CNNs automatically learn spatial hierarchies from radar representations like spectrograms and range-Doppler maps. Kang et al. [24] employed CNNs in a two-step process for segmenting and classifying activity sequences, achieving over 97% segmentation accuracy and 95% classification accuracy with 60 GHz millimeter-wave radar data. Yu et al. [10] introduced a ‘Dual-View CNN’ operating on orthogonal projections of voxelized point clouds, reaching 97.61% accuracy. However, CNNs applied solely on micro-Doppler maps simply treated as images, lacks inherent temporal modeling, a critical limitation for continuous activity recognition.
@@ II-C RNN-based method
To address temporal dynamics in human activities, researchers incorporated Recurrent Neural Networks, particularly Long Short-Term Memory (LSTM[25] and Gated Recurrent Unit (GRU[26] networks, specifically designed for sequential data processing. Shrestha et al. [27] demonstrated the superiority of bidirectional LSTM networks over traditional SVMs and standard LSTMs in continuous activity recognition. Werthen-Brabants et al. [11] proposed an innovative split Bidirectional (Bi)-RNN architecture for real-time fall alerts, using separate forward & backward RNN components for immediate, refined predictions.
@@ II-D Hybrid CNN-RNN-based method
For distributed radar systems, hybrid models achieve an accuracy of 87.1% in continuous activity recognition tasks with only 71 k parameters [14]. Compared to CNN-based models in non-continuous tasks that utilize 135 M parameters [29] and RNN-based models in continuous tasks that utilize 189 M parameters [27] to achieve their best classification accuracy, hybrid models are more lightweight. However, the Floating-point Operation (FLOP) required by the hybrid model [14] is evaluated to be at least 1 G. The arithmetic intensity of CNNs, especially for multi-channel signals, drives their computational complexity.
@@ II-E Transformer and SSM-based method
Recently, researchers have explored transformer-based architectures and SSM in RadHAR. ViT-based methods [16] achieved 92% accuracy with 2.7 M parameters, while [21] reached 92.1% with 769 k parameters on non-continuous FMCW radar data [30], though both demand significant ##\##FLOP per inference. (169 M and 2.41 G, respectively). In continuous HAR, however, ViT appears to under-perform conventional CNN-based models, such as the first to the third blocks of AlexNet and ResNet [17]. SSM-based approaches [19] hit 99.82% accuracy on a non-continuous Continuous Wave (CW) dataset [30] with 8.7 M parameters and 1.22G ##\##FLOP per inference.
III Analysis of Radar Data fed to ViM and Design Motivation
Figure 1: The schematic of conventional Vision Mamba and its three potential issues while classifying the micro-Doppler-based human activity recognition.
To effectively combine the modeling strength of SSM and decrease computational complexity, it is crucial to consider the unique properties of micro-Doppler signatures. To illustrate this, we trace the data flow from input to SSM in Vision Mamba (ViM), highlighting issues specific to radar micro-Doppler-based HAR tasks. Other details will be explained in section IV. As shown in Fig. 1, classical ViM comprises three core components: patch and position embedding, Mamba block with bidirectional branches, and SSMs.
@@ III Analysis of Radar Data fed to ViM and Design Motivation
III-A Issue 1: Micro-Doppler signatures exhibit significant sparsity.
@@
The micro-Doppler signature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ serves as input to the classifier, with the Doppler frequency dimension represented as the image height $H$ and time as the image width $W$. The number of channels $C$ may correspond to channels from multiple radar sensors or one of the RGB channels. Analysis of Dataset [31] CI4R reveals that $\mathbf{X}$ exhibits an average spectrogram sparsity of approximately 87%, as illustrated in Fig.1 Issue 1. In the subsequent Segment step, $\mathbf{X}$ is divided into a matrix of smaller patches $\mathbf{X}_{seg}\in\mathbb{R}^{C\times I\times J\times H_{seg}\times W_{seg}}$, where $H_{seg}$ and $W_{seg}$ denote the height and width of each patch, respectively. Consequently, the number of patches per column is $I=\frac{H}{H_{seg}}$, and per row is $J=\frac{W}{W_{seg}}$. As depicted in Fig.1 Issue 1, many patches in $\mathbf{X}_{seg}$ contain only background information and resemble one another, providing little discriminative detail regarding Doppler frequency and time within each patch.
III-B Issue 2: Destruction of Spatial and Temporal Information.
Changes in Doppler frequency over time are critical for distinguishing various human activities. However, rectangular patch segmentation disrupts the relative positioning of Doppler frequency and time through two processes:
III-B1 Segment Undermines Temporal Interpretability in SSM, as Shown in Fig. 1, Issue 2 Left
As illustrated in Fig.1 left, a patch with column index $j$ can be interpreted as time point $t_{j}$, where $j=0,1,\cdots J$. During segmentation, the patch matrix of dimensions $I\times J$ is reshaped into a one-dimensional sequence $N=I\cdot J$. This reshaping results in a time point sequence of dimension $N$ as $[t_{0},t_{1},\cdots,t_{J},t_{0},t_{1},\cdots,t_{J},\cdots]$, represented by the yellow folded line in Fig.1 left. However, in a real-time sequence, $t_{J}$ and $t_{0}$ are not continuous. When the dimension $N$ remains unchanged from Segment to SSMbranch, and the SSM branches treat $N$ as the direction of temporal evolution, the SSM encounters a discontinuous sequence. This discontinuity hinders its ability to discern significant features in the micro-Doppler signature.
III-B2 Flatten Compromises Spatial Information Integrity Across Time and Doppler Dimensions in Patches, as Shown in Fig. 1, Issue 2 Right
Following segmentation, each patch is flattened into a vector:
$\mathbf{X^{\prime}_{seg}}=\textit{Flatten}(\mathbf{X_{seg}})$ (1)
where $\mathbf{X^{\prime}{seg}}\in\mathbb{R}^{N\times CH{seg}W_{seg}}$. A pixel with row index $h_{seg}$ can be interpreted as Doppler frequency point $f_{h_{seg}}$, where $h_{seg}=0,1,\cdots H_{seg}$. The Flatten operation results in a Doppler point sequence in $H_{seg}W_{seg}$ as $[f_{0},f_{1},\cdots,f_{h_{seg}},f_{0},f_{1},\cdots,f_{h_{seg}},\cdots]$, as shown in Fig. 1 left. However, $f_{h_{seg}}$ and $f_{0}$ are not continuous in the Doppler frequency sequence, disrupting the Doppler frequency features. This disruption also affects the time dimension within each patch.
In summary, the second issue encountered by conventional ViM is that Doppler and temporal information within $H_{seg}$ and $W_{seg}$ is compromised by flattening.
III-C Issue 3: Challenges for Projection Layers in Capturing Doppler Patterns

### v3

# RadMamba: Efficient Human Activity Recognition through a Radar-based Micro-Doppler-Oriented Mamba State-Space Model

@@ Abstract
Radar-based Human Activity Recognition (HAR) is an attractive alternative to wearables and cameras because it preserves privacy, and is contactless and robust to occlusions. However, dominant Convolutional Neural Network (CNN)- and Recurrent Neural Network (RNN)-based solutions are computationally intensive at deployment, and recent lightweight Vision Transformer (ViT) and State Space Model (SSM) variants still exhibit substantial complexity. In this paper, we present RadMamba, a parameter-efficient, micro-Doppler–oriented Mamba SSM tailored to radar HAR under on-sensor compute, latency, and energy constraints typical of distributed radar systems. RadMamba combines (i) channel fusion with downsampling, (ii) Doppler-aligned segmentation that preserves the physical continuity of Doppler over time, and (iii) convolutional token projections that better capture Doppler-span variations, thereby retaining temporal–Doppler structure while reducing the number of Floating-point Operations/Inference (#\#FLOP/Inf.). Evaluated across three datasets with different radars and types of activities, RadMamba matches the prior best $99.8\%$ accuracy of a recent SSM-based model on the Continuous Wave (CW) radar dataset, while requiring only $1/400$ of its parameters. On a dataset of non-continuous activities with Frequency Modulated Continuous Wave (FMCW) radar, RadMamba remains competitive with leading $92.0\%$ results using about $1/10$ of the parameters, and on a continuous FMCW radar dataset it surpasses methods with far more parameters by at least $3\%$, using only $6.7\,\mathrm{k}$ parameters. Code: https://github.com/lab-emi/AIRHAR.
@@ I Introduction
Human Activity Recognition (HAR) technologies are central to healthcare, elderly care, smart homes, and security. Traditional HAR solutions rely on wearable sensors or cameras. While wearables can offer precise monitoring, they suffer from issues related to user compliance, battery life, and comfort. Camera-based approaches raise privacy concerns and are at high risk of degradation under challenging environmental conditions. Consequently, Radar-based Human Activity Recognition (RadHAR) has emerged as a compelling alternative, leveraging active electromagnetic sensing to infer behaviors while preserving privacy, maintaining robustness to illumination and occlusions, and enabling through-wall sensing [1, 2].
Early RadHAR studies progressed from simple, isolated activities [3, 4] to more realistic settings involving continuous activity monitoring [5, 6], multi-modal and distributed sensing [7, 8], and advanced classifiers. Conventional machine learning with hand-crafted micro-Doppler features (e.g., centroid, bandwidth, statistical descriptors, and Singular Value Decomposition (SVD) components) [4, 5, 6, 9] demonstrated proof of concept but struggled with scalability and generalization in continuous scenarios.
Deep learning improved accuracy by learning spatio-temporal representations directly from radar signatures. CNN-based methods [10, 11, 12, 13], RNN-based methods [14, 7], and hybrid CNN-RNN models [15, 16, 17] substantially advanced performance. However, they often have high computational cost and latency, which limits on-sensor deployment in resource-constrained radar nodes. More recently, transformer and SSM-based architectures [18, 19, 20, 21, 22, 23] have shown strong modeling capacity and explored lightweight designs, but balancing parameter count with computational efficiency, measured by the number of floating-point operations per inference (#FLOP/Inf.), remains challenging for mobile and distributed applications such as gesture interfaces [24] and smart-home radar networks [8].
Micro-Doppler signatures encode the evolution of Doppler frequency over time. Preserving this physical continuity is crucial for maintaining the high accuracy of sequential models; however, common vision-style rectangular patching and tokenization can disrupt Doppler–time ordering and waste computational resources on sparse, low-information regions. Moreover, edge radar deployments impose strict limits on memory, latency, and energy per inference. These constraints motivate architectures that (i) respect the physical features of radar signatures and sparsity patterns and (ii) reduce arithmetic intensity without sacrificing discriminative Doppler structure.
In this work, we propose RadMamba, a radar-oriented, parameter-efficient Mamba SSM for micro-Doppler classification. RadMamba replaces vision-style patching and purely linear token projections with three radar-driven mechanisms: (i) channel fusion with downsampling to mix channels and shrink low-information regions; (ii) Doppler-aligned segmentation that preserves Doppler–time continuity for state-space modeling; and (iii) convolutional token projections to capture Doppler-span variations better and improve inter-patch coherence for the SSM backbone.
The main contributions of this work are:
  • A radar-centric state-space architecture that preserves the physical structure of micro-Doppler signatures via Doppler-aligned segmentation and convolutional token projections, reducing #\#FLOP/Inf. while maintaining temporal–Doppler fidelity.
  • A comprehensive evaluation across CW and FMCW datasets covering non-continuous and continuous settings: RadMamba attains $99.8\%$ on a dataset measured by the Defence Institute of Advanced categorization problems and classification Technology, India (DIAT) [25], $91.2\%$ on a dataset from the laboratory of Computational Intelligence for Radar (CI4R) [26] (competitive with heavier models), and $89.3\%$ on a dataset collected at the University of Glasgow in 2020 (UoG20) [27] (at least $3\%$ higher than prior methods with far more parameters), with substantially fewer parameters and lower #\#FLOP/Inf. across all three datasets.
  • An ablation study showing the synergistic effect of (i) channel fusion with downsampling, (ii) Doppler-aligned segmentation, and (iii) convolutional projections, each contributing to efficiency and accuracy beyond conventional ViT/SSM patching and linear projections.
@@
  • An open-source PyTorch framework, RadHAR, to facilitate reproducible research and rapid prototyping for micro-Doppler HAR.
@@ II-A Conventional machine learning method
These methods involved extracting domain-specific features from radar signals, followed by classification using algorithms like k-Nearest Neighbors (kNN) and Support Vector Machine (SVM). For instance, Erol et al. [5] used kNN to analyze power burst curves in spectrograms for fall detection, distinguishing fall from non-fall events. Li et al. [6] applied SVMs to micro-Doppler features such as Doppler centroid, bandwidth, and SVD components, while Ding et al. [9] extracted 28 features of four types: dynamic Doppler frequency, range change, energy change, and dispersion of range and Doppler from dynamic range-Doppler trajectories for subspace kNN classification. Although effective in controlled environments, these approaches struggled to generalize across diverse conditions and activities due to their dependence on hand-crafted features, often leading to overfitting and poor scalability for continuous tasks [1, 2].
@@ II-B CNN-based method
The shift to deep learning marked a turning point for RadHAR, with CNNs leading the charge. Unlike traditional methods, CNNs automatically learn spatial hierarchies from radar representations like spectrograms and range-Doppler maps. Kang et al. [28] employed CNNs in a two-step process for segmenting and classifying activity sequences, achieving over 97% segmentation accuracy and 95% classification accuracy with 60 GHz millimeter-wave radar data. Yu et al. [13] introduced a ‘Dual-View CNN’ operating on orthogonal projections of voxelized point clouds, reaching 97.61% accuracy. However, when CNNs are applied only to micro-Doppler maps treated as images, they lack inherent temporal modeling, which is a critical limitation for continuous activity recognition.
@@ II-C RNN-based method
To address temporal dynamics in human activities, researchers have incorporated RNNs, particularly Long Short-Term Memory (LSTM[29] and Gated Recurrent Unit (GRU[30] networks, which are specifically designed for sequential data processing. Shrestha et al. [27] demonstrated the superiority of bidirectional LSTM networks over traditional SVMs and standard LSTMs in continuous activity recognition. Werthen-Brabants et al. [14] proposed an innovative split Bidirectional (Bi)-RNN architecture for real-time fall alerts, using separate forward & backward RNN components for immediate, refined predictions.
@@ II-D Hybrid CNN-RNN-based method
For distributed radar systems, hybrid models achieve an accuracy of 87.1% in continuous activity recognition tasks with only 71 k parameters [17]. Compared to CNN-based models in non-continuous tasks that utilize 135 M parameters [32] and RNN-based models in continuous tasks that utilize 189 M parameters [27] to achieve their best classification accuracy, hybrid models are more lightweight. However, the #\#FLOP/Inf. required by the hybrid model [17] is evaluated to be at least 1 G. The arithmetic intensity of CNNs, especially for multi-channel signals, drives their computational complexity.
@@ II-E Transformer and SSM-based method
Recently, researchers have explored transformer-based architectures and SSM in RadHAR. ViT-based methods [19] achieved 92% accuracy with 2.7 M parameters, while [33] reached 92.1% with 769 k parameters on non-continuous FMCW radar data [25], though both demand significant #\#FLOP per inference. (169 M and 2.41 G, respectively). In continuous HAR, however, ViT appears to underperform conventional CNN-based models, such as the first three blocks of AlexNet and ResNet [20]. SSM-based approaches [22] hit 99.82% accuracy on a non-continuous CW dataset [25] with 8.7 M parameters and 1.22G #\#FLOP per inference. Compared to ultra-lightweight object detection models optimized for edge devices, such as YOLO-Fastest with 350 k parameters, and 252 M #FLOP/Inf, directly applying these advanced architectures to radar-based HAR tasks remains costly due to their high computational overhead.
III Proposed RadMamba
To effectively combine the modeling power of SSM with lower computational cost, it is essential to exploit the distinctive characteristics of micro-Doppler signatures. However, applying a conventional Vision Mamba (ViM) to RadHAR reveals three key issues:
@@ III Proposed RadMamba
  • Issue 1: Micro-Doppler signatures exhibit significant background sparsity, so most time–frequency bins carry little task-relevant information.
  • Issue 2: Rectangular patch segmentation disrupts the continuous spatio-temporal structure of micro-Doppler patterns, degrading representation quality.
  • Issue 3: Standard linear projection layers struggle to capture Doppler-specific, time–frequency patterns.
To clarify these points, we trace the data flow from input to SSM in RadMamba, highlighting issues specific to radar micro-Doppler-based HAR tasks and describing our proposed solutions in RadMamba. As shown in Fig. 1, RadMamba consists of three core components: preprocessing and segmentation; patch and positional embedding; and a convolution-projection–based Mamba block with bidirectional SSMs. We customize the architecture for micro-Doppler signals using three solutions: channel fusion with downsampling, Doppler-aligned segmentation, and the Convolutional Projection-based Mamba (CP-Mamba) block.
Figure 1: The schematic of the proposed RadMamba architecture and its novel Solution 1 for radar micro-Doppler-based classification.
III-A Solution 1: Channel Fusion and Downsampling
@@
The micro-Doppler signature $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ serves as input to the classifier, with the Doppler frequency dimension represented as the image height $H$ and time as the image width $W$. The channel dimension $C$ corresponds either to multiple radar signal channels or to RGB channels when visualized as an image. In the subsequent Segment step, $\mathbf{X}$ is partitioned into a matrix of smaller patches, denoted as $\mathbf{X}_{seg}$. As depicted in Fig.1 VisionMamba Issue 1, many patches in $\mathbf{X}_{seg}$ contain only background information and therefore appear visually redundant. Based on our analysis of the CI4R Dataset [26], the micro-Doppler spectrograms exhibit an average spectrogram sparsity of approximately 87%. Such sparse background regions contribute limited discriminative information regarding Doppler frequency or temporal dynamics, reducing the representational effectiveness of these patches.
To address the high sparsity inherent in micro-Doppler signatures, we introduce a preprocessing pipeline, Channel Fusion and Downsampling (Chan-DS), before the segmentation stage, as depicted in Fig. 1.
$\mathbf{X_{c-d}=\text{Chan-DS}(\mathbf{X})}$ (1)
where $\mathbf{X}_{c-d}\in\mathbb{R}^{C_{c-d}\times H_{c-d}\times W_{c-d}}$ with $H_{c-d}<H$ and $W_{c-d}<W$ is the micro-Doppler signature after channel fusion and downsampling. The architecture of the Chan-DS block is as follows:
$\text{Chan-DS}=[$ $(\text{Conv2D},\text{BatchNorm})\times L,$
$\text{Maxpool2D},$
$\text{Maxpool1D},$
$\text{Avgpool1D}\text{ (optional)}]$
where $L=1,2$. The Two-Dimensional (2D) convolutional layer and batch normalization layer mix channel information and extract preliminary features from the input tensor $\mathbf{X}$. A 2D max pooling layer and a One-Dimensional (1D) max pooling layer reduce spatial dimensions. Optionally, an additional 1D average pooling layer further downsamples the time dimension if the data sparsity along the time axis is excessive for FMCW radar datasets.
This approach avoids generating numerous empty or redundant patches. Additionally, downsampling reduces the spatial dimensions before segmentation, thereby further contributing to an energy-efficient design.
III-B Solution 2: Doppler-Aligned Segmentation and Patch Embedding
Figure 2: VisionMamba Issue 2 and RadMamba Solution 2: Doppler-Aligned Segmentation
The feature map after segmentation $\mathbf{X}_{seg}$ has dimensions of $C\times I\times J\times H_{seg}\times W_{seg}$, where $H_{seg}$ and $W_{seg}$ denote the height and width of each patch. Thus, the number of patches per column and per row are determined by $I=\frac{H}{H_{seg}}$ and $J=\frac{W}{W_{seg}}$, respectively, resulting in a total number of patches $N=IJ$.
As illustrated in Fig. 2 (a) (upper), patches indexed by $j$ along the horizontal direction correspond to temporal positions $t_{j}$, where $j=0,1,\ldots,J$. Within each patch, as shown in Fig. 2 (a) (bottom), pixels indexed by $h_{seg}$ along the vertical direction correspond to Doppler frequency points $f_{h_{seg}}$, where $h_{seg}=0,1,\ldots,H_{seg}$. Each patch is then flattened into a vector:
$\mathbf{X^{\prime}_{seg}}=\text{Flatten}(\mathbf{X_{seg}})$ (2)
where $\mathbf{X^{\prime}_{seg}}\in\mathbb{R}^{N\times CH_{seg}W_{seg}}$.
The second limitation of conventional VisionMamba lies in the fact that patch flattening disrupts Doppler–temporal continuity, as depicted in Fig. 2 (a). On the one hand, rectangular segmentation results in a temporal ordering in $\mathbf{X}^{\prime}_{seg}$ along the dimension $N$ of the form:
$[t_{0},t_{1},\cdots,t_{J},t_{0},t_{1},\cdots,t_{J},\cdots]$ (3)
represented by the yellow folded arrow in Fig. 2 (a). Since VisionMamba treats $N$ as a temporal evolution dimension in the SSM branches, the artificial jump from $t_{J}$ to $t_{0}$ introduces a temporal discontinuity, making it difficult for the SSM to extract meaningful temporal dynamics. On the other hand, the flattening operation yields a Doppler frequency sequence inside each patch of the form:
$[f_{0},f_{1},\cdots,f_{h_{seg}},f_{0},f_{1},\cdots,f_{h_{seg}},\cdots]$ (4)
as shown in Fig. 2 bottom. The discontinuity between $f_{h_{seg}}$ and $f_{0}$ breaks the frequency pattern within each patch, simultaneously degrading both Doppler frequency consistency and the temporal evolution represented in that patch.
To address this issue, we propose Doppler-aligned segmentation, as shown in Fig. 2 (b):
$\mathbf{X_{seg}}=\text{DA-Seg}(\mathbf{X}_{c-d})$ (5)
This Doppler-aligned segmentation ensures that the temporal evolution of the patch sequence remains strictly ordered as:
$[t_{0},t_{1},\cdots,t_{J}]$ (6)
corresponding directly to the time sequence $W$. When the segmented data is fed into the SSM, as shown in Fig. 2 (b) (upper), the SSM receives a continuous temporal input stream. Furthermore, Doppler-aligned segmentation produces a patch vector with dimension
$CH_{seg}W_{seg}=H_{c-d}$ (7)
preserving the entire Doppler profile within each patch. Consequently, after flattening, the Doppler frequency sequence remains intact as:
$[f_{0},f_{1},\cdots,f_{h\_seg}]$ (8)
as illustrated in Fig. 2(b) (bottom). This maintains Doppler continuity, enabling the model to leverage fine-grained frequency dynamics without disruption.
III-C Solution 3: Convolution Projections in CP-Mamba Blocks